Find the margin of error for the given values of c, s, and n


c=0.90

s=2.9

n=50


1
Expert's answer
2021-11-29T15:00:39-0500

"ME=t_{0.05,n-1}\frac{s}{\sqrt{n}}=1.6766\frac{2.9}{\sqrt{50}}=0.6876."